# On-the-fly Surrogation for Complex Nonlinear Dynamics*

### v2

@@ I Introduction
Over the past century, linear time-invariant (LTI) dynamic models have been extensively used in engineering and industrial applications due to their simplicity and effectiveness [1]. Nonetheless, the increasing complexity and performance demands of modern systems often involve operational regimes where nonlinear dynamics inherent to physical processes play a more critical role, exceeding the capabilities of the LTI framework. Consequently, industrial applications require the construction of high-fidelity models, which are a detailed, precise representation of a physical system that closely replicates its actual behaviour. While these models provide high accuracy, they come at the cost of increased model complexity, in terms of model order, nonlinear dynamic relationships, dynamic coupling and computational load. This complexity often renders high-fidelity models not directly suitable for simulation, system analysis, controller design or real-time implementations.
@@
These challenges manifested in the need for so-called surrogate models, which are low-complexity approximations/representations of all important information about the dynamics of the high-fidelity model for a given utilization objective. Various surrogate modelling frameworks have been developed, such as linear parameter-varying (LPV) systems [2, 3] and Koopman approaches [4, 5] which aim at giving a simplified, linear representation of the nonlinear system with respect to the input-output signal relations. However, these methods trade linearity for conservativeness and scheduling/state order of the resulting surrogate models, often making them difficult to use without further model reduction. Also applicability of LPV and Koopman modelling methods is limited to relatively moderate state order, and conversion to these forms often requires an explicit analytical form of the original model, or the methods are entirely data-based, requiring extensive trajectory-based representation of the system behaviour. Another widely used approach involves various types of neural networks (NN) or Gaussian process (GP) models to learn surrogate models from input-output simulation data [6, 7, 8]. However, these methods demand significant computational resources for the training process, their success largely depends on the generation of sufficiently informative data sets, and they are inherently approximative in their nature. Furthermore, the inclusion of new data requires retraining the obtained models, which often prevents real-time adaptation of these types of surrogates.
@@
Additionally, computation of gradients of complicated nonlinear functions has become an inexpensive operation thanks to the recent development of automatic differentiation methods [9]. Consequently, obtaining system snapshots, i.e., local linear approximations of a nonlinear dynamical system at a specific point in operating space, is currently a cheap and reliable source of system information. Each snapshot contains a rich representation of the local system behaviour, and by collecting snapshots across the expected operating range, it is possible to generate an informative data set of the system behaviour. Despite these advantages, figuring out from these local linearizations an accurate global representation of system behaviour has been found challenging, often considered to be impossible, in the literature and has led to many local approximation methods since the dawn of gain scheduling [10, 11]. Hence, there is no established technique that systematically exploits local system snapshots to construct a surrogate model capable of capturing the entire global nonlinear dynamics. Existing approaches are either heavily approximative, rely on complex model conversion methods, or data-driven approximations that demand extensive computational resources and retraining upon the acquisition of new data. This gap motivates the need for a surrogate modelling framework that employs local linear information to reconstruct the global dynamics and allows the incorporation of new system observations without intensive computation processes.
@@
In this work, we introduce a novel light-weight surrogate modelling approach that leverages system snapshots to construct a surrogate model whose computation is performed “on-the-fly”, i.e., the surrogate model is not assembled a priori, it is only computed at its evaluation, while it is learning the dynamics of the underlying nonlinear system in a global sense. To achieve this, we rely on two key contributions. First, we extend the radial basis functions (RBF)-based multidimensional scattered data interpolation approach [12, 13, 14, 15] to reconstruct the local linearized dynamics of the system from a set of snapshots from the system. Second, based on the second fundamental theorem of calculus (FTC) [16], we establish a one-to-one correspondence between the nonlinear dynamics and their local linearization. By combining these ingredients with numerical integration, a reliable approximation of the global dynamics at a time moment can be computationally efficiently obtained in the form of a surrogate model from the local dynamics. Compared to other surrogate modelling methods, our method offers two key advantages: (i) it provides a surrogate model that is not approximative in its nature, and we prove that it converges to the true nonlinear system representation as the number of system snapshots increases, and (ii) the construction of the surrogate model is computationally inexpensive and can be updated with new data seamlessly, as it is computed on-the-fly without requiring costly retraining or additional optimization steps.
@@
The paper is structured as follows. In Section II, the surrogate modelling problem of nonlinear systems based on a set of local linearizations is formulated. For this problem, the proposed on-the-fly approach is presented in Section III, while in Section IV, the capabilities of the method are shown in simulation studies of a controlled Van der Pol oscillator and a mass-spring-damper system with nonlinear dynamics. Finally, in Section V, the main conclusions on the achieved results are drawn and possible further research directions are discussed.
@@
Notation: The set of real numbers, integers, and positive integers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{Z}^{+}$, respectively. Let $a,b\in\mathbb{Z}^{+}$. The row-wise vectorization of a matrix $M\in\mathbb{R}^{a\times b}$ is given by $\operatorname{vec}(M)\in\mathbb{R}^{1\times ab}$, while its inverse (under row dimension $a$) is ${\operatorname{mat}_{a}}(\operatorname{vec}(M))=M$. Similarly, consider column-wise vectorization as $\operatorname{vec}^{\top}(M)$ and consecutive application of the operation in the form of $\operatorname{vec}^{\top}(M_{1},\ldots,M_{n})$. The gradient of the function $g:\mathbb{R}^{a}\to\mathbb{R}$ and the Jacobian of $g:\mathbb{R}^{a}\to\mathbb{R}^{b}$ w.r.t $z$ are denoted by $\nabla_{\mathrm{z}}(g)$ and $\mathrm{J}_{\mathrm{z}}(g)$, respectively.
@@ II-A System Description
The function $f:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}}$ is assumed to be at least one-time continuously differentiable, i.e., $f\in\mathcal{C}^{k}$ with $k\in\mathbb{Z}^{+}$, and, for any initial condition $x(t_{0})$ and any input trajectory $u(t)$, the solutions of (1) to be forward complete and unique for all $t\geq t_{0}$. Furthermore, let $\mathcal{X}\coloneq\{\underline{x}_{i}\leq x_{i}(t)\leq\overline{x}_{i}\}_{i=1}^{n_{\mathrm{x}}}\in\mathbb{X}$ denote the admissible state space, where $\underline{x}_{i}$ and $\overline{x}_{i}$ are the respective lower and upper bounds of the $i$-th state variable, which can be defined based on the operating range of the system. Likewise, let $\mathcal{U}:=\{\underline{u}_{i}\leq u_{i}(t)\leq\overline{u}_{i}\}_{i=1}^{n_{\mathrm{u}}}\in\mathbb{U}$ and :=$\mathcal{H}:=\{\underline{\eta}_{i}\leq\eta_{i}\leq\overline{\eta}_{i}\}_{i=1}^{n_{\mathrm{\eta}}}\in\mathbb{H}$ denote the admissible input and parameter spaces, respectively. Furthermore, let $z=\mathrm{vec}^{\top}(x,u,\eta)$ and also consider the admissible space of $\mathcal{Z}=\mathcal{X}\times\mathcal{U}\times\mathcal{H}$.
@@ II-B Data collection
Let $\mathcal{L}_{f}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}\times(n_{\mathrm{x}}+n_{\mathrm{u}})}$ be the linearization operator of $f$, i.e., the operator that provides the linearized dynamics of (1) at a point $(x_{\ast},u_{\ast},\eta_{\ast})\in\mathbb{X}\times\mathbb{U}\times\mathbb{H}$:
$\mathcal{L}_{f}(x_{\ast},u_{\ast},\eta_{\ast})\coloneq\begin{bmatrix}\mathrm{J}_{\mathrm{x}}(f)&\mathrm{J}_{\mathrm{u}}(f)\end{bmatrix}(x_{\ast},u_{\ast},\eta_{\ast}).$ (2)
Given a set of local snapshots, i.e., linearizations of (1):
$\mathcal{M}$ :-$\coloneq\{\underbrace{\mathcal{L}_{f}(x^{(i)},u^{(i)},\eta^{(i)})}_{M_{i}}\}_{i=1}^{N},:$ (3)
where $\Theta\coloneq\{z^{(i)}\}_{i=1}^{N}$ with $z^{(i)}=\mathrm{vec}^{\top}(x^{(i)},u^{(i)},\eta^{(i)})$ is a set of $N\in\mathbb{Z}^{+}$ distinct observation points. The collection of available local information about (1) is denoted by
$\mathcal{D}\coloneq\{\Theta,\mathcal{M}\}.$ (4)
@@ III-A From local linearizations to global representation
As a first step to our objective, we establish a key relationship between the nonlinear state-space representation and its linearized dynamics. Building on the ideas in [16][17], and utilizing the FTC [18], we show that the the true nonlinear state equation given in (1) can be recovered completely from its linearizations given by $\mathcal{L}_{f}$ in (2).
@@
$f(x,u,\eta)=f(x_{\ast},u_{\ast},\eta)+F(x,u,\eta)\begin{bmatrix}x-x_{\ast}\\ u-u_{\ast}\end{bmatrix},$ (6)
with
[…]
where $\lambda\in[0,1]$.
@@
For any nonlinear function $g_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $g_{i}\in\mathcal{C}^{1}$, scalar $\lambda\in[0,1]$ and point $z_{*}\in\mathbb{R}^{n}$, define the following auxiliary function $\bar{g}$ as follows
@@
Using the FTC and the identities $\bar{g}_{i}(1)=g_{i}(z)$, $\bar{g}_{i}(0)=g_{i}(z_{\ast})$, we obtain
[…]
Now, applying the chain rule of the integrand in (8), we obtain
$\frac{d\bar{g}_{i}}{d\lambda}(z;\lambda)=\Bigg{[}\Big{(}\nabla_{\mathrm{z}}g_{i}\Big{)}^{\top}\Big{(}z_{*}+\lambda(z-z_{*})\Big{)}\Bigg{]}(z-z_{\ast}).$ (9)
@@
[…]
Now, consider $g(z):=[g_{1}(z)\ \cdots\ g_{m}(z)]^{\top}$. Then, by stacking (10) column-wise for all $i\in[1,\cdots,m]$, $g$ is expressed as
[…]
@@
Moreover, when $f(0,0,\eta)=0$, which corresponds to an equilibrium point in CT, Theorem 1 can be simplified.
@@
$f(x,u,\eta)=F(x,u,\eta)\begin{bmatrix}x\\ u\end{bmatrix}$ (12)
@@
[…]
@@
The results from Corollary 2 allow the exact realization of the true nonlinear function $f$ from the linearization operator $\mathcal{L}_{f}$, provided $f(0,0,\eta)=0$. Note that when $f(0,0,\eta)\neq 0$, but $(0,0)\in\mathcal{X}\times\mathcal{U}$, we can perform a coordinate transformation to ensure this property.
However, application of Corollary 2 has a major drawback: it requires complete knowledge of the linearization of $f$, i.e., the operator $\mathcal{L}_{f}$, over the entire operating space. As in our problem setting, we only have a finite number of local snapshots of the system, we need to approximate $\mathcal{L}_{f}$ from the set of linearizations $\mathcal{D}$. This is what we tackle in the next subsection.

### v3

@@ I Introduction
Over the past century, linear time-invariant (LTI) dynamic models have been extensively used in engineering and industrial applications due to their simplicity and effectiveness [1]. Nonetheless, the increasing complexity and performance demands of modern systems often involve operational regimes where nonlinear dynamics inherent to physical processes play a more critical role, exceeding the capabilities of the LTI framework. Consequently, industrial applications require the construction of high-fidelity models, which are detailed, precise representations of physical systems, closely replicating their actual behaviour. While these models provide high accuracy, they come at the cost of high model complexity in terms of model order, nonlinear dynamic relationships, dynamic coupling and computational load. This complexity often renders high-fidelity models not directly suitable for simulation, system analysis, controller design or real-time implementations.
@@
These challenges manifested in the need for so-called surrogate models, which are low-complexity approximations/representations of all important information about the dynamics of the high-fidelity model for a given utilization objective. Various surrogate modelling frameworks have been developed, such as linear parameter-varying (LPV) systems [2, 3] and Koopman approaches [4, 5] which aim to provide a simplified linear representation of the nonlinear system with respect to the input-output signal relations. LPV methods based on interpolated linearizations, however, cannot represent transient dynamics between local models. More generally, these methods trade linearity for conservativeness and scheduling/state order of the resulting surrogate models, often necessitating further model reduction. Also, conversion methods to LPV and Koopman forms often require either an explicit analytical form of the original model or an extensive trajectory-based representation of the system behaviour. Another widely used approach involves various types of neural networks (NN) or Gaussian process (GP) models to learn surrogate models from input-output simulation data [6, 7, 8]. These methods, however, are computationally demanding, their success largely depends on the generation of sufficiently informative data sets, and are inherently approximative in their nature. Moreover, incorporating new data requires retraining the obtained models, which often prevents real-time adaptation of these types of surrogates.
@@
On the other hand, computation of gradients of complicated nonlinear functions has become an inexpensive operation thanks to the recent development of automatic differentiation methods [9]. Consequently, obtaining system snapshots, i.e., local linear approximations of a nonlinear dynamical system at a specific point in the operating space, is currently a cheap and reliable source of system information. Each snapshot contains a rich representation of the local system behaviour, and by collecting snapshots across the expected operating range, it is possible to generate an informative data set of the system behaviour. Despite these advantages, figuring out from these local linearizations an accurate global representation of the system behaviour has been found challenging, often considered to be impossible in the literature, and has led to many local approximation methods since the dawn of gain scheduling [10, 11]. Hence, there is no established technique that systematically exploits local system snapshots to construct a surrogate model capable of capturing the entire global nonlinear dynamics. Existing approaches are either heavily approximative, rely on complex model conversion methods, or data-driven approximations that demand extensive computational resources and retraining upon the acquisition of new data. This gap motivates the need for a surrogate modelling framework that employs local linear information to reconstruct the global dynamics and allows the incorporation of new system observations without intensive computation processes.
@@
In this work, we introduce a novel light-weight surrogate modelling approach that leverages system snapshots to construct a surrogate model whose computation is performed “on-the-fly”, i.e., the surrogate model is only computed at its evaluation, while it is learning the dynamics of the underlying nonlinear system in a global sense. To achieve this, we rely on two key contributions. First, based on the second fundamental theorem of calculus (FTC) [12], we establish a one-to-one correspondence between the nonlinear dynamics and their local linearization. Second, we extend the radial basis functions (RBF)-based multidimensional scattered data interpolation approach [13, 14] to reconstruct the local linearized dynamics of a system from a finite set of snapshots. By combining these ingredients with numerical integration, a reliable approximation of the global dynamics at a time moment can be computationally efficiently obtained in the form of a surrogate model from the local dynamics. Compared to other surrogate modelling methods, our method offers three key advantages: (i) it provides a surrogate model that is not approximative in its nature, and we prove that it converges to the true nonlinear system representation as the number of local linearizations increases; (ii) the construction of the surrogate model does not require the explicit equations of the underlying system or large data sets but a limited set of local linearizations only, and (iii) can be updated with new data seamlessly without requiring costly retraining or additional optimization steps.
@@
The paper is structured as follows. In Section II, the surrogate modelling problem of nonlinear systems based on a set of local linearizations is formulated. For this problem, the proposed on-the-fly approach is presented in Section III, while in Section IV, the capabilities of the method are shown in simulation studies of a controlled Van der Pol oscillator and a mass-spring-damper system with nonlinear dynamics. Finally, in Section V, the main conclusions on the achieved results and further research directions are discussed.
@@
Notation: The set of real numbers, integers, and positive integers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{Z}^{+}$, respectively. Let $a,b\in\mathbb{Z}^{+}$. The row-wise vectorization of a matrix $M\in\mathbb{R}^{a\times b}$ is given by $\operatorname{vec}(M)\in\mathbb{R}^{1\times ab}$, while its inverse (under row dimension $a$) is ${\operatorname{mat}_{a}}(\operatorname{vec}(M))=M$. Similarly, consider column-wise vectorization as $\operatorname{vec}_{\mathrm{c}}(M)$ and consecutive application of the operation in the form of $\operatorname{vec}_{\mathrm{c}}(M_{1},\ldots,M_{n})$. The gradient of the function $g:\mathbb{R}^{a}\to\mathbb{R}$ and the Jacobian of $g:\mathbb{R}^{a}\to\mathbb{R}^{b}$ w.r.t $z$ are denoted by $\nabla_{\mathrm{z}}(g)$ and $\mathrm{J}_{\mathrm{z}}(g)$, respectively.
@@ II-A System Description
The function $f:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}}$ is assumed to be at least one-time continuously differentiable, i.e., $f\in\mathcal{C}^{k}$ with $k\in\mathbb{Z}^{+}$, and in the CT case, for any initial condition $x(t_{0})$ and any input trajectory $u(t)$, the solutions of (1) to be forward complete and unique for all $t\geq t_{0}$. Furthermore, let $\mathcal{X}\coloneq\{\underline{x}_{i}\leq x_{i}(t)\leq\overline{x}_{i}\}_{i=1}^{n_{\mathrm{x}}}\subseteq\mathbb{X}$ denote the admissible state space, where $\underline{x}_{i}$ and $\overline{x}_{i}$ are the respective lower and upper bounds of the $i$-th state variable, which can be defined based on the operating range of the system. Likewise, let $\mathcal{U}:=\{\underline{u}_{i}\leq u_{i}(t)\leq\overline{u}_{i}\}_{i=1}^{n_{\mathrm{u}}}\subseteq\mathbb{U}$ and $\mathcal{H}:=\{\underline{\eta}_{i}\leq\eta_{i}\leq\overline{\eta}_{i}\}_{i=1}^{n_{\mathrm{\eta}}}\subseteq\mathbb{H}$ denote the admissible input and parameter spaces, respectively. Furthermore, let $z=\operatorname{vec}_{\mathrm{c}}(x,u,\eta)$ and also consider the admissible space of $\mathcal{Z}=\mathcal{X}\times\mathcal{U}\times\mathcal{H}$.
@@ II-B Data collection
Let $\mathcal{L}_{f}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{u}}}\times\mathbb{R}^{n_{\mathrm{\eta}}}\to\mathbb{R}^{n_{\mathrm{x}}\times(n_{\mathrm{x}}+n_{\mathrm{u}})}$ be the linearization operator of $f$, i.e., the operator that provides the linearized dynamics of (1) at a point $(x_{\ast},u_{\ast},\eta_{\ast})\!\in\!\mathcal{X}\!\times\mathcal{U}\!\times\mathcal{H}$:
$\vskip-3.0pt\mathcal{L}_{f}(x_{\ast},u_{\ast},\eta_{\ast})\coloneq\begin{bmatrix}\mathrm{J}_{\mathrm{x}}(f)&\mathrm{J}_{\mathrm{u}}(f)\end{bmatrix}(x_{\ast},u_{\ast},\eta_{\ast}).$ (2)
Consider a set of local snapshots, i.e., linearizations of (1):
$\vskip-2.0pt\mathcal{M}\coloneq\{\underbrace{\mathcal{L}_{f}(x^{(i)},u^{(i)},\eta^{(i)})}_{M_{i}}\}_{i=1}^{N},$ (3)
where $\Theta\coloneq\{z^{(i)}\}_{i=1}^{N}$ with $z^{(i)}=\operatorname{vec}_{\mathrm{c}}(x^{(i)},u^{(i)},\eta^{(i)})$ is a set of $N\in\mathbb{Z}^{+}$ distinct observation points. The collection of available local information about (1) is denoted by
$\vskip-3.0pt\mathcal{D}\coloneq\{\Theta,\mathcal{M}\}.$ (4)
@@ III-A From local linearizations to global representation
As a first step towards our objective, we establish a key relationship between the nonlinear state-space representation and its linearized dynamics. Building on the ideas in [15][16], and utilizing the FTC, we show that the true nonlinear state equation given in (1) can be recovered completely from its linearizations given by $\mathcal{L}_{f}$ in (2).
@@
$\vskip-2.0ptf(x,u,\eta)=f(x_{\ast},u_{\ast},\eta)+F(x,u,\eta)\begin{bmatrix}x-x_{\ast}\\ u-u_{\ast}\end{bmatrix}$, (6)
where
$F(x,u,\eta)\hskip-1.0pt$ $=\hskip-3.5pt\int\limits_{0}^{1}\hskip-3.5pt\mathcal{L}_{f}\left(x_{\ast}+\lambda(x-x_{\ast}),u_{\ast}+\lambda(u-u_{\ast}),\eta\right)\,\mathop{}\!\mathrm{d}\lambda.$
@@
For any nonlinear function $g_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $g_{i}\in\mathcal{C}^{1}$, scalar $\lambda\in[0,1]$ and point $z_{*}\in\mathbb{R}^{n}$, we define the following auxiliary function $\bar{g}$ as follows
@@
By the FTC and the identities $\bar{g}_{i}(z;1)=g_{i}(z)$, $\bar{g}_{i}(z;0)=g_{i}(z_{\ast})$:
$g_{i}(z)-g_{i}(z_{\ast})=\int\limits_{0}^{1}\frac{\mathop{}\!\mathrm{d}\bar{g}_{i}}{\mathop{}\!\mathrm{d}\lambda}(z;\lambda)\,\mathop{}\!\mathrm{d}\lambda.$ (8)
Now, applying the chain rule of the integrand in (8):
$\frac{\mathop{}\!\mathrm{d}\bar{g}_{i}}{\mathop{}\!\mathrm{d}\lambda}(z;\lambda)=\Bigg{[}\Big{(}\nabla_{\mathrm{z}}g_{i}\Big{)}^{\top}\Big{(}z_{*}+\lambda(z-z_{*})\Big{)}\Bigg{]}(z-z_{\ast}).$ (9)
@@
$g_{i}(z)=g_{i}(z_{\ast})+\int\limits_{0}^{1}\Bigg{[}\Big{(}\nabla_{\mathrm{z}}g_{i}\Big{)}^{\top}\Big{(}z_{*}+\lambda(z-z_{*})\Big{)}\Bigg{]}\,\mathop{}\!\mathrm{d}\lambda\,(z-z_{\ast}).$ (10)
Then, consider $g(z):=[g_{1}(z)\ \cdots\ g_{m}(z)]^{\top}$. By stacking (10) column-wise for all $i\in[1,\cdots,m]$, $g$ is expressed as
$g(z)=g(z_{\ast})+\int_{0}^{1}\mathrm{J}_{\mathrm{z}}\Big{(}g(z_{\ast}+\lambda(z-z_{\ast}))\Big{)}\,\mathop{}\!\mathrm{d}\lambda\,(z-z_{\ast}).$ (11)
@@
When $f(0,0,\eta)=0$, which corresponds to an equilibrium point in CT, Theorem 1 can be simplified.
@@
$\vskip-3.0ptf(x,u,\eta)=F(x,u,\eta)\begin{bmatrix}x\\ u\end{bmatrix}$, (12)
@@
$F(x,u,\eta)$ $=\int_{0}^{1}\mathcal{L}_{f}\left(\lambda x,\lambda u,\eta\right)\,\mathop{}\!\mathrm{d}\lambda.$
@@
The results from Corollary 2 allow the exact realization of the true nonlinear function $f$ from the linearization operator $\mathcal{L}_{f}$, provided $f(0,0,\eta)=0$. Note that when $f(0,0,\eta)\neq 0$, but $(0,0)\in\mathcal{X}\times\mathcal{U}$, we can perform a coordinate transformation to ensure this property. However, application of Corollary 2 has a major drawback: it requires complete knowledge of the linearization of $f$, i.e., the operator $\mathcal{L}_{f}$, over the entire operating space. As in our problem setting we only have a finite number of local snapshots of the system, we need to approximate $\mathcal{L}_{f}$ from the set of linearizations $\mathcal{D}$. This is what we tackle in the next subsection.